INFRARED SPECTROSCOPY OF THE MASS 31 CATION: PROTONATED FORMALDEHYDE VS. THE TRIPLET METHOXY CATION

\underline{J. D. MOSLEY}, T. C. CHENG, and M. A. DUNCAN, University of Georgia, Dept. of Chemistry, 1001 Cedar St, Athens, GA 30602.

The m/z=31 cation is produced by ionization and fragmentation of methanol, ethanol, dimethyl ether, etc. Two structures have been proposed, protonated formaldehyde (^{1}CH_{2}OH^{+}) and the triplet methoxy cation (^{3}CH_{3}O^{+}). The infrared spectrum of the mass 31 cation is obtained using infrared photodissociation spectroscopy with Ar tagging. The spectrum reveals the presence of two stable isomers, protonated formaldehyde (^{1}CH_{2}OH^{+}) and the triplet methoxy cation (^{3}CH_{3}O^{+}). The triplet methoxy cation has been studied extensively and is predicted to interconvert to protonated formaldehyde through an essentially barrierless process on a timescale much faster than our experiment (> 100 μs). The presence of two structural isomers is verified by comparison of spectra from different precursors and spectra of different temperature ions from the same precursor.